solve the linear system by gaussian elimination

 x − y + 2z − w = −1,

2x + y − 2z − 2w = −2 ,

−x + 2y − 4z + w = 1 ,

3x − 3w = −3


1
Expert's answer
2022-02-18T01:43:53-0500

Augmented matrix


"\begin{bmatrix}
 1 & -1 & 2 & -1 & & -1 \\
 2 & 1 & -2 & -2 & & -2\\
 -1 & 2 & -4 & 1 & & 1 \\
 3 & 0 & 0 & -3 & & -3 \\
\end{bmatrix}"

"R_2=R_2-2R_1"


"\begin{bmatrix}
 1 & -1 & 2 & -1 & & -1 \\
 0 & 3 & -6 & 0 & & 0\\
 -1 & 2 & -4 & 1 & & 1 \\
 3 & 0 & 0 & -3 & & -3 \\
\end{bmatrix}"

"R_3=R_3+R_1"


"\begin{bmatrix}
 1 & -1 & 2 & -1 & & -1 \\
 0 & 3 & -6 & 0 & & 0\\
 0 & 1 & -2 & 0 & & 0 \\
 3 & 0 & 0 & -3 & & -3 \\
\end{bmatrix}"

"R_4=R_4-3R_1"


"\begin{bmatrix}
 1 & -1 & 2 & -1 & & -1 \\
 0 & 3 & -6 & 0 & & 0\\
 0 & 1 & -2 & 0 & & 0 \\
 0 & 3 & -6 & 0 & & 0 \\
\end{bmatrix}"

"R_2=R_2\/3"


"\begin{bmatrix}
 1 & -1 & 2 & -1 & & -1 \\
 0 & 1 & -2 & 0 & & 0\\
 0 & 1 & -2 & 0 & & 0 \\
 0 & 3 & -6 & 0 & & 0 \\
\end{bmatrix}"

"R_1=R_1+R_2"


"\begin{bmatrix}
 1 &0 & 0 & -1 & & -1 \\
 0 & 1 & -2 & 0 & & 0\\
 0 & 1 & -2 & 0 & & 0 \\
 0 & 3 & -6 & 0 & & 0 \\
\end{bmatrix}"

"R_3=R_3-R_2"


"\begin{bmatrix}
 1 &0 & 0 & -1 & & -1 \\
 0 & 1 & -2 & 0 & & 0\\
 0 & 0 & 0 & 0 & & 0 \\
 0 & 3 & -6 & 0 & & 0 \\
\end{bmatrix}"

"R_4=R_4-3R_2"


"\begin{bmatrix}
 1 &0 & 0 & -1 & & -1 \\
 0 & 1 & -2 & 0 & & 0\\
 0 & 0 & 0 & 0 & & 0 \\
 0 & 0 &0 & 0 & & 0 \\
\end{bmatrix}"

If "w=t, t\in \R, z=s, s\in \R," then "x=-1+t, y=2s, z=s, w=t, t,s\in \R."


The linear system has infinitely many solutions


"(-1+t, 2s, s, t), \ t,s\in \R"